• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 13:41:12 on July30,2005

    Entry number 150105

    This entry is a followup to: 150102

    keyword=Septa reset procedure note

    We discover empirically:
    after the septum IOC hangs and has to be reset (can be done
    via hareboot), and after the IOC comes back to life and one gets screens such 
    as shown in 
    halog entry 150102 
    that you need to click the Power Supply CMD "ON" button to get 
    the temperatures to reset; then the Interlock Reset will clear 
    the Interlocks (assuming the temperatures are OK). 
    Moral: Don't just wait for the yellow crossed-out lines
    on the temperature readouts to clear themselves, try the power
    on button...